A leading magazine (like Barrons) reported at one time that the average number of weeks an individual is unemployed s 30 weeks. Assume that for the population of all unemployed individuals the population mean length of unemployment is 30 weeks and that the population standard deviation is 4 weeks. Suppose you would like to select a random sample of 34 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is less than 31 Pix<31)- Find the probability that a sample of size n-34 is randomly selected with a mean less than 3I PIM<31)- Enter your answers as numbers accurate to 4 decimal places. Points possible: 1 Submilt Next 1/23/2019 BANG &OLUFSE 5 6 Right answer needed please.
Answer #1

Let's consider X be the length of unemployment in weeks.

Given mean of X μ-30

standard deviation is sigma=4

n=34

a) The probability that the single randomly selected value is less than 31 is given by

P(X<31)=P(rac{{X}-mu}{{sigma }}<rac{{31}-mu}{{sigma }})

31 - 30

P(X < 31) = P(Z <-)

P(X < 31) = P(Z < 0.25)

P(X < 31) 0.5987 ### By using z table

b) The probability that the sample of size 34 is randomly selected with mean less than 31 is given by

P(M<31)=P(rac{M-mu}{rac{sigma }{sqrt{n}}}<rac{31-mu}{rac{sigma }{sqrt{n}}})

31 - 30 34

P(M<31)=P(Z<rac{1}{rac{4 }{5.83}})

P(M<31)=P(Z<rac{1}{0.6861})

P(M<31)=P(Z<1.46)

P(M < 31) 0.9279 ### By using the z table.

The probability that the sample of size 34 is randomly selected with mean less than 31 =0.9279